OpenStudy (anonymous):

find f(1) of 5x(cosx-sinx)+(sinx+cosx)(5)

OpenStudy (anonymous):

10cos(1)

OpenStudy (anonymous):

cos(1) as in cos(2pi)?

OpenStudy (anonymous):

5*x*(Cos[x] - Sin[x]) + (Sin[x] + Cos[x])*5 = 5 (1 + x) Cos[x] = 10 Cos[1] = 5.403